### v2

# Active Sparse Bayesian Committee Machine Potential for Isothermal-Isobaric Molecular Dynamics Simulations

Abstract
[…]
I Introduction
In the last two decades, the application of machine learning potentials (MLPs) in the field of materials simulations has proven to be highly successful, particularly for the accurate prediction of potential energy surfaces (PESs) of the reference ab initio calculations,Unke et al. (2021); Deringer et al. (2021); Bartók et al. (2010); Chen et al. (2023); Gilmer et al. (2017); Batzner et al. (2022); Hajibabaei et al. (2021); Hajibabaei and Kim (2021); Hajibabaei et al. (2021); Myung et al. (2022); Vandermause et al. (2022, 2020); Metcalf et al. (2020); Na and Kim (2021); Zhung et al. (2024); Imbalzano et al. (2021) such as density functional theory (DFT),Laasonen et al. (1993); Richards et al. (2016); Boero et al. (1998) many-body perturbation (MP),Willow et al. (2015); Del Ben et al. (2013, 2015) and coupled-cluster (CC) theories. Moreover, MLPs have an affordable computational cost, allowing one to investigate previously inaccessible chemical phenomena at the first-principles level with currently accessible supercomputers. This breakthrough has paved the way for large-scale and long-time simulations in various scientific domains, including physics, chemistry, biology, and materials science. Several distinct MLP methods have been developed, such as the Behler-Parrinello MLP,Behler and Parrinello (2007); Behler (2011); Pun et al. (2019); Eckhoff and Behler (2021); Schran et al. (2020, 2021); Kapil et al. (2022) Gaussian approximation potentials (GAPs),Bartók et al. (2010, 2013); Bartók and Csányi (2015); Klawohn et al. (2023) sparse Gaussian process regression (SGPR).Hajibabaei et al. (2021); Hajibabaei and Kim (2021); Quiñonero-Candela and Rasmussen (2005); Hajibabaei et al. (2021); Myung et al. (2022), graph neural network potentials (GNNs),Gilmer et al. (2017); Batzner et al. (2022) gradient-domain machine learning (GDML),Chmiela et al. (2017) moment tensor potentials (MTP),Shapeev (2016); Podryabinkin and Shapeev (2017); Novikov et al. (2021) and equivariant message-passing neural networks.Batatia et al. (2022); Batzner et al. (2022); Batatia et al. (2022) The versatility of these methodologies has led to their widespread adoption and successful application across different scientific disciplines, promising valuable insights and outcomes.Lim and Jung (2021)
The Gaussian Process ($\mathcal{GP}$) based MLP boasts in the active learning, uncertainty prediction, and low data requirement.Rasmussen and Williams (2005); Shahriari et al. (2016); Lawrence (2005); Álvarez and Lawrence (2011) However, despite all these advantages, the computational scalability of $\mathcal{GP}$-based MLPs comes under scrutiny. $\mathcal{GP}$-based MLPs face computational challenges when dealing with a large dataset, whose training complexity increases dramatically to $\mathcal{O}(n^{3})$ with the size of kernel matrix $n$ due to its inversion. As a result, $\mathcal{GP}$s are restricted to handling the datasets whose size are smaller than $10^{4}$. This limitation led to the development of sparsification techniques to select a subset of data point $m$ among $n$ data points to reduce the computational complexity.Hajibabaei et al. (2021); Hajibabaei and Kim (2021); Quiñonero-Candela and Rasmussen (2005); Hajibabaei et al. (2021); Myung et al. (2022) However, the sparsification approaches including SGPR still struggles with large number of inducing sets $m$ as $\mathcal{O}(nm^{2})$ for training. Another strategy to mitigate the steep computational scaling is to use aggregation models, including the Bayesian Committee Machine (BCM)Tresp (2000). The BCM, for instance, combines the predictions from $P$ distinct local expert models, each trained on different sub-data. This further reduces the computational scaling to $\mathcal{O}(nm^{2}/P^{3})$ for training.
MD simulations serve as a powerful tool to complement experimental findings by offering a detailed view of the physical system at the atomic scale. By comparing the results of MD simulations with their corresponding experimental observations (i.e., phase diagrams, reaction kinetics, and structure factor), one can effectively test the accuracy and reliability of the given PES. To perform isobaric simulations, such as isothermal-isobaric ($NpT$) and isobaric-isenthalpic ($NpH$) MD simulations, it is crucial to accurately estimate the pressure of system. In this regard, improving the accuracy of pressure estimation of MLPs could enhance the reliability of isobaric ML MD simulations. While other types of MLPs such as neural network MLP Wang et al. (2018) and GAPs,Klawohn et al. (2023) incorporated the estimation of virial pressure, SGPR-based MLP has not yet included the estimation of virial pressure.Hajibabaei et al. (2021)
[…]
Figure 1: Schematic of active sparse Bayesian committee machine (BCM) potential. (a) The entire dataset is divided into $P$ distinct subsets. Within each subset, a sparse Gaussian process regression (SGPR) MLP is trained. Subsequently, a combined version of MLP for the entire dataset is built through the BCM algorithm. The weights $\frac{\beta_{\alpha}}{\sigma_{\alpha}^{2}}$ ensure that the predictions from confident models are more outspoken than those from less confident models. (b) Illustration of active BCM algorithm. Each local expert SGPR model undergoes training up to a predefined kernel size. Before surpassing the kernel size limit, our algorithm initiates training for the next SGPR model. This process is sequentially executed, resulting in the assembly of $P$ local SGPR MLPs, BCM(1:$P$), forming a Bayesian committee. Throughout the MD simulation, this ensemble of $P$ committee MLPs predicts the energy, forces, and virial tensor of the system.
II Methods
All MD simulations were carried out using the ASE package.Larsen et al. (2017) We employed a time step of 0.5 fs for the integration of the equations of motion that were propagated according to combined Nosé-Hoover and Parrinello-Rahman dynamics.Melchionna et al. (1993); Melchionna (2000); Holian et al. (1990) To acquire the training and testing datasets for analysis, three different systems were chosen. For liquid boron nitride (BN) and LGPS, we trained the BCM potential to the density functional theory (DFT) at the Perdew-Burke-Ernzerhof (PBE) level.Perdew et al. (1996) All the DFT calculations with projector augmented wave (PAW) pseudopotentialsKresse and Joubert (1999) were performed using the Vienna Ab initio Simulation Package (VASP).Kresse and Furthmüller (1996, 1996) We performed isothermal ($NVT$) MD simulations of liquid boron nitride (BN) at temperature of $T=$ 7500 K. The liquid BN system consists of 64 atoms within the cubic box of 7.252 Å. We performed isothermal-isobaric ($NpT$) MD simulations of LGPS at the ambient pressure and temperature ranging from $T=$ 600 to 1000 K. A $\Gamma$-point sampling was employed with an energy cutoff set at 500 eV. We sampled various structures of ice phases (II, III, and V) and liquid water through $NpT$ MD simulations using the MB-pol force fieldBabin et al. (2013, 2014); Riera et al. (2023) at pressures of $p=$ 0.4 and 0.6 GPa, with a temperature of $T=$ 255 K.
III Theory
III.1 Virial Kernel in Sparse Gaussian Process Regression potential
The SGPR energy is given as decomposed local energy function of a local chemical environment (LCE) $\rho_{i}$ as
$E=\sum_{i=1}^{N}\mathcal{E}(\rho_{i}),$ (1)
where $\mathcal{E}(\rho_{i})$ is a local energy of LCE $\rho_{i}$. This indicates that the SGPR energy as well as force acting on atom $i$ depends on the relative coordinates of surrounding atoms because the LCE of atom $i$ is defined by
$\rho_{i}=\{\vec{r}_{ij};j\neq i\quad|\vec{r}_{ij}|<r_{c}\},$ (2)
where $\vec{r}_{ij}=\vec{r}_{j}-\vec{r}_{i}$, $r_{c}$ is a cutoff radius for neighborhood relations.
The core component of SGPR is building a covariance kernel $\mathcal{K}(\rho_{i},\rho_{j})$ that encodes the similarity between $\rho_{i}$ and $\rho_{j}$. The kernel is invariant with respect to the symmetry operations that leaves the potential energy invariant, e.g. translations and rotations. Here, the smooth overlap of atomic positions (SOAP)Bartók et al. (2013) is used for the descriptor of a similarity kernel between LCEs and is defined as
[…]
where $\hat{\bm{R}}$ is the three-dimensional rotation operator, $\xi_{i}(\vec{r})$ is the atomic density in the neighborhood of $i$,
$\xi_{i}(\vec{r})=\sum_{j\in\rho_{i}}e^{\alpha|\vec{r}-\vec{r}_{ij}|^{2}},$ (4)
and $\alpha$ is a hyperparameter.
In the SGPR potential, we select a reduced set of LCEs $z=\{\chi_{j}\}_{j=1}^{m}$, called inducing set, which are significantly distinct from each other and are sufficient statistics for the original data $X=\{R_{1},\cdots,R_{n}\}$ to improve the scalability. It is likely that many LCE pairs in $X$ are similar: $\mathcal{K}(\rho_{i},\rho_{j})\approx 1$. With the inducing set of LCEs, the latent local energy of atom $i$ is expressed as
$\mathcal{E}(\rho_{i})=\sum_{j=1}^{m}\mathcal{K}(\rho_{i},\chi_{j})~{}w_{j}$ (5)
where $\bm{w}=[w_{1},\cdots,w_{m}]$ are the trainable weights for the descriptors in $z$. The potential energy for a given configuration $R$ becomes
$E(R)=\sum_{i}^{N}\mathcal{E}(\rho_{i})=\sum_{i,j}\mathcal{K}(\rho_{i},\chi_{j})~{}w_{j}.$ (6)
The forces $f_{i}^{\mu}$ are obtained from
$f_{i}^{\mu}=\sum_{j=1}^{m}\left[-\sum_{k=1}^{N}\frac{\partial\mathcal{K}(\rho_{k},\chi_{j})}{\partial r_{i}^{\mu}}\right]w_{j}=\sum_{j=1}^{m}\dot{\mathcal{K}}_{i}^{\mu}(R,\chi_{j})~{}w_{j},$ (7)
where $\dot{\mathcal{K}}_{i}^{\mu}(R,\chi_{j})$ is the covariance kernel for the force on atom $i$ in $\mu$-direction, $\mu=\{x,y,z\}$.
The previous SGPR potential didn’t consider the pressure tensor in the training process.Hajibabaei et al. (2021) Due to this, the SGPR potential often couldn’t accurately predict the pressure tensor of a given system. To improve this shortcoming, our strategy is to incorporate the pressure data into the training process of SGPR. The instantaneous pressure tensor $\mathcal{P}^{\mu\nu}$ is defined as
$\mathcal{P}^{\mu\nu}=Nk_{B}\mathcal{T}^{\mu\nu}/V+\mathcal{W}^{\mu\nu}/V,$ (8)
where $\mathcal{T}^{\mu\nu}$ is the instantaneous temperature tensor. To predict the pressure tensor, we build the internal virial tensor $\mathcal{W}^{\mu\nu}$ as below
$\mathcal{W}^{\mu\nu}$ $=$ $\sum_{i}^{N}f_{i}^{\mu}r_{i}^{\nu}$ (9)
$=$ $\sum_{i}^{N}\left[\sum_{j}^{m}\dot{\mathcal{K}}_{i}^{\mu}(R,\chi_{j})~{}w_{j}\right]r_{i}^{\nu}.$
The weight vector $\bm{w}$ is obtained by the fitting of $E(R)$, $f_{k}^{\mu}$, and $\mathcal{W}^{\mu\nu}$ to the target energies, forces, virial pressures $\bm{y}_{n}$ for a given dataset $X=\{R_{1},\cdots,R_{n}\}$. The subsequent fitting process of SGPR is nicely reformulated as a form of linear equation with the regularization hyperparameter $\sigma$, Hajibabaei et al. (2021)
$\begin{bmatrix}\bm{K}_{nm}\\ \sigma\bm{L}^{T}\end{bmatrix}\bm{w}_{m}=\begin{bmatrix}\bm{y}_{n}\\ \bm{0}\end{bmatrix},$ (10)
where $\bm{L}$ is the Cholesky factor of $\bm{K}_{mm}$ ($\bm{K}_{mm}=\bm{L}\bm{L}^{T}$, where $\bm{L}$ is lower triangular) and $\bm{0}$ is a columnar vector of zeros with length $m$.
III.2 Active Sparse Bayesian Committee Machine potential
The SGPR algorithm still encounters difficulties in obtaining the solution $\bm{w}_{m}$ of Eq. (10), which involves the inversion of a kernel matrix, with large size of training datasets $X$ and inducing set $z$ of LCEs. In order to mitigate this challenge, we employed BCMTresp (2000) to combine SGPR MLPs which are separately trained on different subsets as a local expert (Figure 1 and S1). When building MLP of ice polymorphs, rather than utilizing all ice phases as a training dataset, we partitioned each ice phase into training subsets. Subsequently, we trained each local expert models for corresponding subset of ice phase. We combined these individually trained local expert SGPR models into a universal BCM MLPs. This approach not only streamlines the training process but also addresses the computational challenges associated with the inversion of large kernel matrices, paving a way for building an efficient and accurate universal MLPs.
For given subsets of the training dataset $\{n_{1},\cdots,n_{P}\}$ and inducing LCEs $\{m_{1},\cdots,m_{P}\}$, the BCMTresp (2000) potential energy is
$E=\hat{S}\sum_{\alpha=1}^{P}\frac{\beta_{\alpha}}{\sigma_{\alpha}^{2}}E_{\alpha},$ (11)
where the potential energy $E_{\alpha}$ of $\alpha$-th SGPR local expert ($R_{n\in\alpha}$ and $z_{\alpha}=\{\chi_{j}\}_{j\in\alpha})$ is given
$E_{\alpha}(R)$ $=$ $\sum_{i,(j\in m_{\alpha})}\mathcal{K}(\rho_{i},\chi_{j})~{}w_{j}^{\alpha}.$ (12)
where $\bm{w}^{\alpha}$ are the training weights of $\alpha$-th local SGPR model.
The forces $f_{i}^{\mu}$ and internal virial $\mathcal{W}^{\mu\nu}$ of the BCM potential are obtained from
$f_{i}^{\mu}$ $=$ $\hat{S}\sum_{\alpha=1}^{P}\frac{\beta_{\alpha}}{\sigma_{\alpha}^{2}}f_{i,\alpha}^{\mu}$ (13)
$\mathcal{W}^{\mu\nu}$ $=$ $\hat{S}\sum_{\alpha=1}^{P}\frac{\beta_{\alpha}}{\sigma_{\alpha}^{2}}\mathcal{W}^{\mu\nu}_{\alpha},$ (14)
where $f_{i,\alpha}^{\mu}$ and $\mathcal{W}^{\mu\nu}_{\alpha}$ represent the forces and the internal virial of the $\alpha$-th SGPR local expert, which are given
$f_{i,\alpha}^{\mu}$ $=$ $\sum_{j\in m_{\alpha}}\dot{\mathcal{K}}_{i}^{\mu}(R,\chi_{j})~{}w_{j}^{\alpha},~{}~{}\mathrm{and}$
$\mathcal{W}^{\mu\nu}_{\alpha}$ $=$ $\sum_{i}^{N}f_{i,\alpha}^{\mu}r_{i}^{\nu}.$
The maximum of the covariance loss for the $\alpha$-th subset LCEs, denoted as $\sigma_{\alpha}^{2}$, is used to weight the $\alpha$-th committee prediction. The $\sigma_{\alpha}^{2}$ is estimated by
$\sigma_{\alpha}^{2}=\mathrm{max}(1-\bm{K}_{*m}\bm{K}_{mm}^{-1}\bm{K}_{*m}^{T}),$ (15)
where $\bm{K}_{*m}$ is the covariance matrix between a test configuration $\bm{x}^{*}$ and $\alpha$-th inducing set $z_{\alpha}$. Furthermore, we weight the prediction with $\beta_{\alpha}$ following the concepts of the robust BCM where the individual committee prediction is weighted by the differential entropy, $\beta_{\alpha}=-\log(\sigma_{\alpha}^{2})$. In the BCM, a particular local expert which has the highest similarity with a test configuration $\bm{x}^{*}$ mostly contributes to the energy, forces, and pressure predictions. The final normalization factor $\hat{S}$ is thus given:
$\hat{S}$ $=$ $\left[\sum_{\alpha}^{P}\frac{\beta_{\alpha}}{\sigma_{\alpha}^{2}}\right]^{-1}.$ (16)
The significant benefit of employing the SGPR-based BCM is that, in addition to the sparsification, we invert an approximate $(m/P)$-dimensional matrix instead of inverting an $m$-dimensional matrix ($m=\sum_{\alpha}^{P}m_{\alpha}$). Hence, the computational cost of the matrix inversion scales as $\mathcal{O}(m^{3}/P^{3})$ rather than $\mathcal{O}(m^{3})$. Compared to “product of experts” (POE),Hajibabaei and Kim (2021) in which the potential energy prediction is $E\approx\frac{1}{P}\sum_{\alpha}^{P}E_{\alpha}$, the BCM scale factor $\beta_{\alpha}/\sigma_{\alpha}^{2}$ ensures that larger weights are applied for the predictions of confident models than less confident models.
IV Results and Discussion
IV.1 Accurate pressure predictions of virial kernel method
In Section III.1, the weight vector $\bm{w}_{m}$ relies on the covariance matrix $\bm{K}_{nm}$ and corresponding target values for energy, forces, virial pressure denoted as $\bm{y}_{n}$. The weight vector’s values vary based on whether the covariance matrix and corresponding target values include the virial pressure term. In this subsection, we compare instantaneous pressure predictions using the weights obtained under two different approaches. This comparative analysis provides insights into the effect of virial kernel in covariance matrix and into the accuracy of pressure prediction. In our subsequent discussion, we will elucidate how the omission of the virial kernel, $\mathcal{W}^{\mu\nu}$, can precipitate abrupt transitions in energy or pressure during the update in active learning. Such discontinuities are problematic in $NpT$ or $NpH$ simulations, undermining its reliability. This phenomenon could raise significant concerns regarding the credibility of simulation results in the absence of virial kernel in SGPR potential.
The first scheme (Scheme 1) involves building $\bm{w}_{m}$ solely with $K_{nm}$ and $\bm{y}_{n}$ that includes energies and forces. The second scheme (Scheme 2) trains $\bm{w}_{m}$ by incorporating virial kernel into $\bm{K}_{nm}$ and $\bm{y}_{n}$ includes energies, forces, and pressures. To substantiate the predictive power of the virial kernel, we evaluated the accuracy of Scheme 1 and Scheme 2 across different phases of materials: high pressure liquid phase BN at $T=$ 7500 K ($NVT$) (Figure S2), solid Li electrolyte LGPS at temperatures ranging from $T=$ 600 to 1000 K ($NpT$), and coexisting phase of ice III and liquid water at pressures of $p=$ 0.4 and 0.6 GPa, with a temperature of $T=$ 255 K ($NpH$). Through these systematic investigations across different systems and phases, we aim to offer a comprehensive evaluation of the virial kernel’s predictive accuracy.
Figure 2: Comparison of pressure predictions with and without virial kernel. Three different systems were tested to assess the impact of including the virial term in the covariance matrix $\bm{K}_{nm}$ on pressure calculations: (a) liquid boron nitride, (b) solid Li electrolyte LGPS, and (c) a coexisting system of ice III and liquid water. Scheme 1 (red dots) involves constructing $\bm{w}_{m}$ solely with $K_{nm}$ and target values of energies and forces. Scheme 2 (blue dots) incorporates virial kernel into $\bm{K}_{nm}$ and $\bm{y}_{n}$ includes target energies, forces, and pressure to obtain $\bm{w}_{m}$. Under high pressure conditions, the predictive accuracy of Scheme 1 for pressure exhibits significant errors.
Once we trained the weight vector $\bm{w}_{m}$, we estimated the pressures with the test dataset using Eqs. (8) and (9). This highlights that Scheme 2 is capable of accurately reproducing target instantaneous pressures as opposed to Scheme 1 (Figure 2). BN is an industrially important material which is manufactured at high temperature and pressure. Recently, it was proposed that amorphous BN exhibits low dielectric characteristics and can be used for high performance electronic applications.Hong et al. (2020) Under extreme thermodynamic conditions (T = 7500 K), Scheme 1 struggles to predict the pressure of liquid BN system, whose error is substantially large with the RMSE of 38.43 GPa. However, the inclusion of virial kernel improves its prediction from the RMSE of 38.43 GPa to 6.54 GPa (Figure 2a). The LGPS is one of solid state Li ion electrolytes, and at room temperature, Li ions diffuse while remaining backbone maintains its crystallinity (Figure S3-5). Although the Li ions diffuse away, as it maintains the solid phase, the predictions of both Scheme 1 and Scheme 2 closely align with the reference values with the average predicted pressure value of approximately $p=$ 0 GPa (Figure 2b). Although there was a slight improvement in pressure prediction from the RMSE of 0.26 GPa to 0.13 GPa, the integration of the virial term into the covariance matrix did not manifest a significant effect when the system is in its solid phase under ambient pressure. Under the mild conditions (0.2 GPa $\leq$ $p$leq 0.6 GPa) for the coexisting ice-liquid water system, we again see the pressure prediction improvement by incorporating the virial kernel from RMSE of 1.30 GPa to 0.048 GPa (Figure 2c). Based on the above observations, we conclude that it is essential to incorporate the virial kernel function into covariance matrix, particularly when dealing with a system under high pressure.
IV.2 On-the-flying active learning $NpT$ MD simulations of Liquid Water
On-the-flying active learning $NpT$ MD simulations were performed on liquid water under a pressure of $p=0.2$ GPa and a temperature of $T=270$ K, subsequent to training an expert model through active learning $NVT$ MD simulations at the same temperature. The initial liquid system consisted of 256 water molecules within a cubic box measuring 19.715 Å in box length. Due to its inability to precisely estimate pressure, the $NpT$ MD simulation without the inclusion of virial kernel (Scheme 1) failed to adjust the box’s fluctuations. Updating the dataset $X$, inducing set $z$, and weight vector $\bm{w}$ led to a noticeable sudden change in pressure, disrupting the isenthalipic and isobaric conditions of MD simulations (see Figure 3). On the other hands, when the virial kernel in included in the covariance matrix (Scheme 2), both energy and pressure predictions are smooth during the update in on-the-flying active learning $NpT$ MD simulation, resulting in precise adjustment of the box. In this study, hence, we trained SGPR local expert MLP models for ice II, ice III, and ice V through active learning MD simulations under constant pressure and constant temperature conditions, using an active sparse BCM potential that includes the virial kernel (Scheme 2).
Figure 3: Fluctuations of potential energy $E_{\mathrm{pot}}$, pressure $p_{z}$, and cell $L_{z}$ with and without virial kernel during the active learning $NpT$ MD simulation. Scheme 1 shows abrupt discontinuity in energy (upper) and pressure (middle) according to updating dataset, inducing set, and weight vector. As a result, the $NpT$ barostat with Scheme 1 fails for box adjustment leading to a sudden box shrinkage (lower).
Figure 4: Ice-water phase diagram. Ice-water melting points are calculated through $NpH$ MD simulations of coexisting ice-liquid systems using the MB-pol (circle lines) and BCM (star dashed lines). Solid black line represents the experimental melting line.Petrenko and Whitworth (2002)
IV.3 Coexistence line of ice-liquid water
Anisotropic $NpH$ MD simulations were employed to compute the melting temperature ($T_{m}$) of ice within the coexisting ice-liquid phase. $NpH$ MD simulations enable the spontaneous adjustment of temperature to satisfy the condition $\mu_{\mathrm{ice}}(p,T)_{T=T_{m}}=\mu_{\mathrm{liq}}(p,T)_{T=T_{m}}$. Specifically, when the temperature of ice-liquid coexisting system is higher than $T_{m}$, the chemical potential of the ice phase is higher than that of the liquid phase: $\mu_{\mathrm{ice}}(p,T)_{T>T_{m}}>\mu_{\mathrm{liq}}(p,T)_{T>T_{m}}$. Consequently, the ice phase undergoes a process of melting with the absorption of heat, leading to a reduction in the kinetic temperature of the coexisting system towards the melting temperature $T_{m}$. One notable advantage of employing anisotropic $NpH$ MD simulations over $NVE$ MD simulation is the absence of the pressure-anisotropy problem, as the three principle components of the pressure tensor can be adjusted to the specified external pressure in the former (see Figure S7).Yoo et al. (2004, 2009, 2009); Vega et al. (2008)
We estimate the melting temperatures of ice II, III, and V under pressure conditions of 0.2 $\sim$ 0.6 GPa using the MB-pol water modelRiera et al. (2023). To achieve this, on-the-fly active learning SGPR $NpT$ MD simulations were performed to create BCM MLP for each coexisting systems: ice II-water, ice III-water, and ice V-water. Three separatedly trained SGPR local expert MLP models were then combined into one sparse robust BCM potential. This combined potential was used for $NpH$ MD simulations of coexisting ice-liquid systems. Non-active $NpH$ MD simulations were performed to determine the melting temperatures of ice II, III, and V. Throughout these MD simulations, we maintained the conservation of enthalpy ($H$), while observing fluctuations in the total energy, which is the sum of the potential energy and kinetic energy (see Figure S8).
Figure 4 illustrates the melting points of ice II, III, and V as determined from our $NpH$ MD simulations using the MB-pol water model. The ice-water phase diagram predicted with the MB-pol potential closely align with those obtained from enhanced-coexistence simulations, as reported by Bore and PaesaniBore and Paesani (2023) (see Figure S10). The BCM MLP stabilizes ice V over ice II, consistent with Zhang’s predicted phase diagram.Zhang et al. (2021) The Bore and PaesaniBore and Paesani (2023) using a sophisticated deep neural network potential, however, observed that ice V became a metastable phase, absent from the phase diagram. We attribute this discrepancy to the failure in incorporating high model uncertainty configurations of the water-ice system in the training dataset. In contrast, our active BCM potential adeptly learns model uncertainties on-the-fly, enhancing model accuracy and thereby stabilize ice V in the phase diagram (Figure 4). However, the melting points predicted by the BCM MLP exhibit a slight upward shift in temperature relative to the benchmark MB-pol potential (Figure 4). This deviation primarily stems from the noise present in the energy and force predictions of the BCM potential. Notwithstanding the shifts in melting temperatures, we emphasize the BCM potential’s capability to accurately replicate the physical and thermodynamic properties of the reference MB-pol water model. While the discrepancies in melting temperatures suggest a room for further improvements in accuracy of model, the significance lies in the BCM potential’s reproduction of reference phase diagram. This affirms the BCM potential’s utility in capturing the essence of water’s behavior under various conditions while reducing the computational scaling of training SGPR MLP for large data.
Deep neural network ML modelsZhang et al. (2021); Bore and Paesani (2023) based on DFT and MB-pol predict that ice III is metastable when compared to ice II. Our BCM MLP based on MB-pol exhibits the same trend, indicating that ice III is indeed metastable in comparison to ice II. The MB-pol water model, used in MD simulations, lacks the ability to allow proton transfer within ice, limiting its capacity to represent all possible configurations of proton-disordered ice. Consequently, the entropy stemming from proton disorder is not factored into the chemical potential of ice $\mu_{\mathrm{ice}}(p,T)$. This discrepancy extends to $NpH$ MD simulations of ice-liquid systems, where the entropy attributed to proton disorder remains unaddressed. The disparity between these models and experimental data likely arises from neglecting the entropy arising from proton disorder in ice III.MacDowell et al. (2004); Vega et al. (2008) Notably, the entropy of proton-disordered ice III surpasses that of proton-ordered ice II. Correctly incorporating the entropy related to proton disorder in ice III should lead to an expectation that the melting temperature of proton-disordered ice III exceeds that of proton-ordered ice II.MacDowell et al. (2004)
Table 1 summarizes the computational costs of obtaining the weight vector $\bm{w}_{m}$ in Eq. (10). Each SGPR local expert for ice II-water, ice III-water, and ice V-water has a comparable number of descriptors in the inducing set $z$. Consequently, the computational cost for acquiring $\bm{w}_{m}$ increases proportionally with the total number of descriptors in the dataset $\mathcal{O}(n_{d})$. Combining all SGPR local experts into a single SGPR model results in a slightly increase in both $m$ and $n_{d}$ compared to the SGPR local expert for ice II-water. Nevertheless, the computational cost in obtaining $\bm{w}_{m}$ nearly doubles in comparison with the SGPR local expert of ice II-water, satisfying that the computational cost scales as $\mathcal{O}(n_{d}m^{2})$.
Table 1:  The computational cost of obtaining the solution $\bm{w}_{m}$ in Eq. (10). $n$, $n_{d}$ and $m$ represent the number of data in the dataset, the total number of descriptors in the dataset, and the number of descriptors in the inducing set $z$.
SGPR local expert $n$ $n_{d}$ $m$ CPU time (s)
Ice V 13 26208 261 2.17
Ice III 21 40824 300 4.28
Ice II 25 64800 284 6.69
Ice II+Ice III+Ice V 32 78624 357 11.73
V Conclusions
In this study, we introduced two theoretical advancements to enhance the precision of pressure predictions and to reduce the computational scaling of SGPR MLPs. For the isobaric MD simulations (i.e. $NpT$ and $NpH$), accurate predictions of system pressure is crucial. We showed that the introduction of a virial kernel into the covariance matrix substantially improves pressure predictions, particularly for liquid systems under extreme conditions. This improvement is important for the fidelity of isobaric MD simulations, particularly for the study of phase transition and the behavior under extreme conditions. Furthermore, the refined MLP with the virial kernel enables the estimation of the melting temperature of ice polymorphs. We achieve this by employing $NpH$ MD simulations of the ice-water coexisting phases for three cases (ice II, ice III, and ice V).
We also developed the active BCM potential, which addresses the computational scalability challenges inherent to $\mathcal{GP}$-based potentials, including SGPR. By aggregating local sparse SGPR kernel potentials into a unified BCM model, the active BCM potential effectively reduces the dimensionality of the kernel matrix, thus reducing the computational overhead without the need for training a large dataset. This approach not only streamlines the model training process but also enhances the model’s transferability to a broader range of systems and conditions. Importantly, the active BCM potential facilitates on-the-fly learning of model uncertainties, further refining the model’s accuracy and reliability.
These two approaches—the virial kernel and the active BCM potential—mark a significant step forward in the application of SGPR potential to isobaric MD simulations. By addressing key challenges in pressure estimation and computational efficiency, our study paves the way for more accurate, reliable, and scalable MD simulations across a wide array of materials and conditions. These advancements hold promise for deepening our understanding of material behaviors and phase transitions, contributing valuable insights to the fields of chemistry, physics, and materials science.
Author Contributions
Soohaeng Yoo Willow: Conceptualization, Methodology, Software, Validation, Formal analysis, Data Curation, Writing, Visualization, Funding acquisition. Dong Geon Kim: Data Curation, Writing, Visualization. R. Sundheep: Data Curation, Writing. Amir Hajibabaei: Conceptualization, Software. Chang Woo Myung: Conceptualization, Methodology, Writing, Visualization, Formal analysis, Supervision, Project administration, Funding acquisition.
Conflicts of interest
There are no conflicts to declare.
Acknowledgements
We are grateful to V. Kapil, S. Pourasad, and C. Schran for helpful discussions. SYW and CWM acknowledge the support from the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00222245). CWM acknowledges the support from the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1C1C1010605). The authors are grateful for the computational support from the Korea Institute of Science and Technology Information (KISTI) for the Nurion cluster (KSC-2021-CRE-0542, KSC-2022-CRE-0115, KSC-2022-CRE-0424, KSC-2023-CRE-0059, KSC-2023-CRE-0198, KSC-2023-CRE-0332, KSC-2023-CRE-0355, KSC-2023-CRE-0454).
Data availability
See the supplementary material for a detailed compilation of the obtained results as well as further data and analysis to support the points made throughout the text. The input and output files associated with this study and all analysis can be found on GitHub.